# Disentangling Brillouin’s negentropy law of information and Landauer’s law on data erasure

### v2

@@ IV.3 Pair of states versus path
According to Landauer “Logical irreversibility is associated with physical [thermodynamical] irreversibility”[8]. This is a claim of generality of Landauer’s law and is probably the most problematic point.
Logical irreversibility refers to a loss of information (hence Landauer’s idea of establishing a link with data erasure) we possess about a system. Loss of information that is an increase in the quantity of information we lack to describe the system, or an increase of the uncertainty about it; in other words, an increase in its entropy. Logical irreversibility is simply a positive change in entropy. It is solely related to a change in a state quantity. This is a consequence of the Clausius definition of entropy as a state quantity and of the successive mathematical results of Gibbs and Shannon equating quantity of information and entropy, all condensed in Brillouin’s law of information.
@@
How does Landauer arrive at the conclusion that the two irreversibilities are associated? In fact, Landauer forces the path to go through a particular non-differentiable sequence of events, so that a property which initially in the theory is related solely to a peculiar path (thermodynamic irreversibility), becomes also a property of an ordered pair of states (logical irreversibility). This viewpoint either empties the very definition of a state of all meaning, or transforms information and thus entropy into quantities which are not state quantities. In both cases, Landauer’s idea is incompatible with thermodynamics and information theory and leads to inconsistencies.
Consider a system changing from A to B, with $S_{\text{A}}<S_{\text{B}}$. The change is logically irreversible (the information we have decreases), thus (according to Landauer), it is also always thermodynamically irreversible. But the change in the other direction from B to A is logically reversible, thus (according to Landauer), it could be thermodynamically reversible. In summary: the change from A to B would be thermodynamically irreversible, and that from B to A would be thermodynamically reversible. What exactly would “thermodynamically reversible” mean in this context?
In fact, Landauer erasure and Brillouin’s law only lead to the same result in the context of a cyclic data acquisition (such as that of a Szilard machine). Outside of this context, they do not. Brillouin’s law concerns the acquisition of information, Landauer’s law concerns the erasure of data, hence the contradiction if we equate the two. In the context of a Szilard engine and Landauer’s scenario, the data erased before acquisition relates to the engine’s state during the previous cycle. This data is essentially obsolete; it concerns an information about the past. The erased data no longer contains an information about the engine’s current state.
@@ V Conclusive epistemological point
In the literature, Landauer’s law on data erasure is generally presented as a decisive improvement to information theory that provides us with the missing key to understanding the link between information and thermodynamics. For instance, one can read: “The Landauer principle is one of the cornerstone of the modern theory of information” (Herrera [12]).
@@
This situation is very strange because there is absolutely no experiment, no observation, that can be explained by Landauer’s law and that would not be explained without it. On one side, the link between information and energy, as established by Maxwell or Szilard engines, is perfectly explained by Brillouin’s law. On the other side, the erasure of data-bits, whatever their form, is also perfectly explained by thermodynamics. Landauer’s law adds absolutely nothing to the theory, except inconsistencies that make it incompatible with thermodynamics and information theory. So, how do we conceive its popularity? It is as if an explanation that consists solely of deducing observations (past and future) from a set of theoretical statements (principles and laws) is insufficient. Yet, that is precisely the only thing expected of theory.
@@
The popularity of Landauer’s law can only be due to a misconception of what are concepts in physics. There is no physical theory without concepts. But concepts such as energy and information (but also that of field and many others) are not material entities, although they are fully physical, in the sense that they play a crucial role in current physical theories. They are, in fact, conventions. The meaning of any concept, what it encompasses, its definition, is a convention because we cannot see it and consequently, establish a direct link between what we see and what we mean. We cannot show a concept to someone.
@@
“Information is physical” (Landauer [9]). “Information is not a disembodied abstract entity” (Landauer [10]). Landauer’s law on data erasure is in reality an attempt to materialize information, which, without Landauer, would remain an abstract concept. According to his idea, the smallest piece of information is actually localized and materialized under the form of a data-bit. In this sense, Landauer’s idea belongs to the reductionist branch of materialism. It is in direct continuity with the initial intention of Gibbs’ statistical mechanics. Here, the purpose is neither to contest materialism, nor to deny the utility of reductionism when it allows an economy of thought. It is simply to deny reductionism as a profession of faith or a creed.
@@
The concept of energy is widely used in physics although it is a pure concept. There is no such thing as an “energy particle” (an object) that could exist independently of us. Energy is like a universal. And nobody has a problem with that. Why not consider information in the same way?
@@
Attached to materialism and to “the scientific conception of the world”[13] (neo-positivism) is the idea that the only two sources of knowledge are that underlying mathematics, with no other synthetic a priori knowledge, and that of observations. The former imposes a strict requirement of consistency in the statements of the theory, while the latter obliges us to take into account the role of the observer, since there is no observation without observer.
@@
This is where a potential problem of misunderstanding lies. Accounting for the observer introduces information. A concept which is emergent since the observer is only meaningful at the macroscopic scale. Hence the danger of falling into a form of holism that would consist of considering emergence as synthetic a priori knowledge. In other words, precisely what neo-positivism intends to reject. This apparent contradiction introduced by emergence into materialism probably explains the search for a reductionist solution like Landauer’s.
@@
Actually, a theoretical statement involving an emergent property, such as the principle of maximum entropy, is not a synthetic a priori knowledge, it is a convention. A convention that is exactly of the same nature as that of the first principle of thermodynamics or Euclidian geometry [18, 16]. The choice of one convention or another is guided solely by its convenience and in particular by the economy of thought [14] it provides: the simplest is the best. The prior probability distribution of phases of a system, which is directly deduced from the principle of maximum entropy, is in this sense also a convention (it is directly derived from a convention). It does not pretend to reflect the reality of the system. It is not the system that maximizes the probability distribution of its phases, but our conventional representation of the system that must do so if we want to be rational.
@@
Emergentism is generally and historically divided into two categories [53]: ontological and epistemological. The former considers it is inherent in the essence of things, the latter that it is only imposed to us by our necessary limited knowledge. However, with regard to experience, both lead to the same result. The difference is a metaphysical question out of the scope of science. The neo-positivist position [13, 16], inspired by Poincaré’s conventionalism [18], which is defended here, makes it possible to avoid this problem and constitutes a third way which could be called “conventional emergentism” and is not so far removed from pragmatism [54, 55].

### v3

@@ IV.3 Pair of States Versus Path
According to Landauer, “Logical irreversibility is associated with physical [thermodynamical] irreversibility” [37]. This is a claim of generality of Landauer’s law and is probably the most problematic point.
Logical irreversibility refers to a loss of information (hence Landauer’s idea of establishing a link with data erasure) that we possess about a system. A loss of information is an increase in the quantity of information we lack to describe the system, or an increase in the uncertainty about it; in other words, an increase in its entropy. Logical irreversibility is simply a positive change in entropy. It is solely related to a change in a state quantity. This is a consequence of the Clausius definition of entropy as a state quantity and of the successive mathematical results of Gibbs and Shannon equating quantity of information and entropy, all condensed in Brillouin’s law of information.
@@
How does Landauer arrive at the conclusion that the two irreversibilities are associated? In fact, Landauer forces the path to go through a particular non-differentiable sequence of events, so that a property which initially in the theory is related solely to a peculiar path (thermodynamic irreversibility), becomes also a property of an ordered pair of states (logical irreversibility). This viewpoint either empties the very definition of a state of all meaning or transforms information and thus entropy into quantities that are not state quantities. In both cases, Landauer’s idea is incompatible with thermodynamics and information theory and leads to inconsistencies.
Consider a system changing from A to B, with $S_{\text{A}}<S_{\text{B}}$. The change is logically irreversible (the information we lack increases); thus, according to Landauer, it is also always thermodynamically irreversible. But the change in the other direction from B to A is logically reversible; thus, according to Landauer, it could be thermodynamically reversible. In summary, the change from A to B would be thermodynamically irreversible, and that from B to A would be thermodynamically reversible. What exactly would “thermodynamically reversible” mean in this context?
In fact, Landauer erasure and Brillouin’s law only lead to the same result in the context of a cyclic data acquisition (such as that of a Szilard engine). Outside of this context, they do not. Brillouin’s law concerns the acquisition of information; Landauer’s law concerns the erasure of data; hence, the contradiction if we equate the two. In the context of a Szilard engine and Landauer’s scenario, the data erased before acquisition relates to the engine’s state during the previous cycle. This data is essentially obsolete; it concerns information about the past. The erased data no longer contains information about the engine’s current state.
@@ V Conclusive Epistemological Point
In the literature, Landauer’s law on data erasure is generally presented as a decisive improvement to information theory that provides us with the missing key to understanding the link between information and thermodynamics. For instance, one can read: “The Landauer principle is one of the cornerstones of the modern theory of information” (Herrera [25]).
@@
This situation is very strange because there is absolutely no experiment and no observation, which can be explained by Landauer’s law, and that would not be explained without it. On one side, the link between information and energy, as established by the Maxwell or Szilard engine, is perfectly explained by Brillouin’s law. On the other side, the erasure of data bits, whatever their form, is also perfectly explained by thermodynamics. Landauer’s law adds absolutely nothing to the theory, except for inconsistencies that make it incompatible with thermodynamics and information theory. So, how do we conceive its popularity? It is as if an explanation that consists solely of deducing observations (past and future) from a set of theoretical statements (principles and laws) is insufficient. Yet, that is precisely the only thing expected of theory.
@@
The popularity of Landauer’s law can only be due to a misconception of what concepts are in physics. There is no physical theory without concepts. But concepts such as energy and information (but also that of field and many others) are not material entities, although they are fully physical, in the sense that they play a crucial role in current physical theories. They are, in fact, conventions. The meaning of any concept, what it encompasses, and its definition, is a convention because we cannot see it, and consequently, establish a direct link between what we see and what we mean. We cannot show a concept to someone.
@@
“Information is physical” (Landauer [38]). “Information is not a disembodied abstract entity” (Landauer [39]). Landauer’s law on data erasure is, in reality, an attempt to materialize information, which, without Landauer, would remain an abstract concept. According to his idea, the smallest piece of information is actually localized and materialized under the form of a data bit. In this sense, Landauer’s idea belongs to the reductionist branch of materialism. It is in direct continuity with the initial intention of Gibbs’ statistical mechanics. Here, the purpose is neither to contest materialism nor to deny the utility of reductionism when it allows an economy of thought. It is simply to deny reductionism as a profession of faith or a creed.
@@
The concept of energy is widely used in physics, although it is a pure concept. There is no such thing as an “energy particle” (an object) that could exist independently of us. Energy is like a universal and nobody has a problem with that. Why not consider information in the same way?
@@
Attached to materialism and to “the scientific conception of the world” [21] (neo-positivism) is the idea that the only two sources of knowledge are that underlying mathematics, with no other synthetic a priori knowledge, and that of observations. The former imposes a strict requirement of consistency in the statements of the theory, while the latter obliges us to take into account the role of the observer, since there is no observation without an observer.
@@
This is where a potential problem of misunderstanding lies. Accounting for the observer introduces information. A concept that is emergent since the observer is only meaningful at the macroscopic scale. Hence, the danger of falling into a form of holism that would consist of considering emergence as synthetic a priori knowledge. In other words, precisely what neo-positivism intends to reject. This apparent contradiction introduced by emergence into materialism probably explains the search for a reductionist solution like Landauer’s.
@@
Actually, a theoretical statement involving an emergent property, such as the principle of maximum entropy, is not a synthetic a priori knowledge; it is a convention. A convention that is exactly of the same nature as that of the first principle of thermodynamics or Euclidean geometry [48, 8]. The choice of one convention or another is guided solely by its convenience and, in particular, by the economy of thought [41] it provides: the simplest is the best. The prior probability distribution of phases of a system, which is directly deduced from the principle of maximum entropy, is in this sense also a convention (it is directly derived from a convention). It does not pretend to reflect the reality of the system. It is not the system that maximizes the probability distribution of its phases, but our conventional representation of the system that must do so if we want to be rational.
@@
Emergentism is generally and historically divided into two categories [4]: ontological and epistemological. The former considers it to be inherent in the essence of things, whereas the latter holds that it is imposed only on us by our necessary limited knowledge. However, with regard to experience, both lead to the same result. The difference is a metaphysical question out of the scope of science. The neo-positivist position [21, 8], inspired by Poincaré’s conventionalism [48], which is defended here, makes it possible to avoid this problem and constitutes a third way that could be called “conventional emergentism” and is not so far removed from pragmatism [1, 15].